# The CASE Framework - A New Architecture for Participatory Research and Digital Health Surveillance

### v1

@@ 1. Introduction
Digital survey platforms have become essential tools in participatory research with various different tools available across different domains (Wright, 2017). Many provide survey logic capabilities, typically allowing researchers to define conditional branching or skipping questions based on respondent answers. Many platforms also offer dynamic content features where responses can be piped into later questions, allowing basic personalization of the questionnaire experience. Despite this, most mainstream survey platforms still follow a mostly linear, predefined flow. The branching logic is usually limited to conditions based on previous responses within the survey itself.
Truly context-aware or externally adaptive workflows remain rare. Recent analyses highlight that popular survey services ”lack usable mechanisms for seamlessly importing participants’ data from other systems” (Velykoivanenko et al., 2024).
Although several platforms can preload known information or trigger minor follow-ups, there is limited support for dynamically incorporating external context, such as real-time sensor data, environmental conditions, or personal digital data into the survey logic during execution.
@@ 2.2. Related Work
Several participatory surveillance platforms have demonstrated the value of crowd-sourced symptom reporting for the monitoring of infectious diseases. In the US, Flu Near You and its successor Outbreaks Near Me (Boston Children’s Hospital, HealthMap, Flu Lab, and Ending Pandemics, 2025) allowed tens of thousands of volunteers to report influenza-like and COVID-19 symptoms weekly, although their static technical architecture limits dynamic survey adaptation, remains largely undocumented and was not designed for broader reuse (Smolinski et al., 2015). Australia’s FluTracking (Carlson et al., 2019) has similarly shown high public engagement and epidemiological value during seasonal influenza outbreaks, but, like other single-purpose platforms, lacks the technical flexibility to dynamically adapt survey flows or rapidly adapt to new diseases during emerging health threats. The UK’s ZOE COVID Symptom Study demonstrated the scalability and research potential of participatory symptom tracking at unprecedented scale, engaging millions of users and generating findings that directly influenced public health policy, although it remained tailored to COVID-19 and was not designed for adaptation by external research groups (Menni et al., 2020). Many European platforms within the Influenzanet network have since transitioned to the open-source CASE framework, establishing a reusable foundation for multi-study participatory research. This shift reflects a broader trend toward scalable, transparent, and adaptable infrastructures for participatory surveillance, identified as a critical need across global systems (McNeil et al., 2022).
@@
The review of existing platforms reveals several key gaps that CASE addresses. Traditional survey platforms like LimeSurvey and Qualtrics, while user-friendly, lack real-time adaptiveness and external context integration. Mobile data collection tools such as ODK and REDCap focus on form-based data capture but are not able to dynamically adjust to external events once deployed. Context-aware research prototypes demonstrate promising concepts, but remain largely a proof-of-concept. Mobile health research platforms such as ResearchKit are sophisticated, but limited to specific ecosystems and health domains. Furthermore, many commercial platforms have restrictive pricing models and limited customization options for research use cases (Wright, 2017; Sue and Ritter, 2012). CASE contributes to this landscape by implementing an event-driven architecture that enables dynamic survey workflows based on external context, participant state, and temporal factors. The framework has been deployed in several participatory surveillance systems and longitudinal studies. Although primarily applied in health research contexts, its modular design allows adaptation to other domains, as illustrated by its use in analysis of political sentiment during live events. However, like other research frameworks, CASE still requires technical expertise for deployment and customization to specific research needs.
@@ 5.2. Selected Use Cases
The Dutch platform tekenradar.nl enables public participation in tracking tick bites and associated health effects for tick borne disease monitoring. The system has gathered tens of thousands of tick bite reports and supports epidemiological research on Lyme disease and related conditions through flexible survey and notification modules tailored to the experiences reported by individual participants (Garcia-Marti et al., 2018; Hofhuis et al., 2017). Recently rebuilt with the CASE framework, it allows users to report tick encounters and symptoms while supporting follow-up engagement through a multitrack longitudinal study structure.
The Post-COVID Research Portal at postcovidonderzoek.nl serves as a national entry point for individuals with long-term symptoms after COVID-19 infection to participate in ongoing scientific studies (National Institute for Public Health and the Environment , Netherland(2025a), RIVM). Developed from the start with the reworked version of the CASE software, the platform collects baseline and follow-up data while enabling centralized participant recruitment across multiple coordinated research projects. By streamlining intake and periodic follow-up engagement, the system improves long-term cohort management and enables targeted substudies. The platform also supports integration with clinical studies, as participants are invited through the portal to provide blood samples for a biomarker study (National Institute for Public Health and the Environment , Netherland(2025b), RIVM), illustrating how CASE allows the digital research infrastructure to connect efficiently with laboratory workflows.

### v2

@@ 1. Introduction
Digital survey platforms have become essential tools in participatory research with various different tools available across different domains (Wright, 2017). Many provide survey logic capabilities, typically allowing researchers to define conditional branching or skipping questions based on respondent answers. Many platforms also offer dynamic content features where responses can be piped into later questions, allowing basic personalization of the questionnaire experience. Despite this, most mainstream survey platforms still follow a mostly linear, predefined flow. The branching logic is usually limited to conditions based on previous responses within the survey itself. Truly context-aware or externally adaptive workflows remain rare. Recent analyses highlight that popular survey services ”lack usable mechanisms for seamlessly importing participants’ data from other systems” (Velykoivanenko et al., 2024). Although several platforms can preload known information or trigger minor follow-ups, there is limited support for dynamically incorporating external context, such as real-time sensor data, environmental conditions, or personal digital data into the survey logic during execution.
@@ 2.2. Related Work
Several participatory surveillance platforms have demonstrated the value of crowd-sourced symptom reporting for the monitoring of infectious diseases. In the US, Flu Near You and its successor Outbreaks Near Me (Boston Children’s Hospital, HealthMap, Flu Lab, and Ending Pandemics, 2025) allowed tens of thousands of volunteers to report influenza-like and COVID-19 symptoms weekly, although their static technical architecture limits dynamic survey adaptation, remains largely undocumented and was not designed for broader reuse (Smolinski et al., 2015). Australia’s FluTracking (Carlson et al., 2019) has similarly shown high public engagement and epidemiological value during seasonal influenza outbreaks, but, like other single-purpose platforms, lacks the technical flexibility to dynamically adapt survey flows or rapidly adapt to new diseases during emerging health threats. The UK’s ZOE COVID Symptom Study demonstrated the scalability and research potential of participatory symptom tracking at unprecedented scale, engaging millions of users and generating findings that directly influenced public health policy (Menni et al., 2020). Beyond its initial role in tracking COVID-19 symptoms, the ZOE app has evolved into a broader research platform now known as the ZOE Health Study. It engages participants in longitudinal studies of diet, chronic symptoms, and post-COVID-19 syndrome (Rjoob et al., 2025; Bermingham et al., 2024). This illustrates the potential for participatory research platforms to remain relevant beyond acute outbreaks. CASE supports a reusable and adaptive architecture that facilitates such transitions, enabling research continuity across domains. Many European platforms within the Influenzanet network have since transitioned to the open-source CASE framework, establishing a reusable foundation for multi-study participatory research. This shift reflects a broader trend toward scalable, transparent, and adaptable infrastructures for participatory surveillance, identified as a critical need across global systems (McNeil et al., 2022).
@@
The review of existing platforms reveals several key gaps that CASE addresses. Traditional survey platforms like LimeSurvey and Qualtrics, while user-friendly, lack real-time adaptiveness and external context integration. Mobile data collection tools such as ODK and REDCap focus on form-based data capture but are not able to dynamically adjust to external events once deployed. Context-aware research prototypes demonstrate promising concepts, but remain largely a proof-of-concept. Mobile health research platforms such as ResearchKit are sophisticated, but limited to specific ecosystems and health domains. Furthermore, many commercial platforms have restrictive pricing models and limited customization options for research use cases (Wright, 2017; Sue and Ritter, 2012). CASE contributes to this landscape by implementing an event-driven architecture that enables dynamic survey workflows based on external context, participant state, and temporal factors. The framework has been deployed in several participatory surveillance systems and longitudinal studies. Although primarily applied in health research contexts, its modular design allows adaptation to other domains, as illustrated by its use in analysis of political sentiment during live events. This flexibility positions CASE not only as a platform for health-related studies, but also as a foundation for participatory research more generally. However, like other research frameworks, CASE still requires technical expertise for deployment and customization to specific research needs.
@@ 5.2. Selected Use Cases
The Dutch platform tekenradar.nl allows public participation in tracking tick bites and associated health effects for the monitoring of tick-borne diseases. The system has gathered tens of thousands of tick bite reports and supports epidemiological research on Lyme disease and related conditions through flexible survey and notification modules tailored to the experiences reported by individual participants (Garcia-Marti et al., 2018; Hofhuis et al., 2017). Recently rebuilt with the CASE framework, it allows users to report tick encounters and symptoms while supporting follow-up engagement through a multitrack longitudinal study structure.
The Post-COVID Research Portal at postcovidonderzoek.nl serves as a national entry point for individuals with long-term symptoms after COVID-19 infection to participate in ongoing scientific studies (National Institute for Public Health and the Environment , Netherland(2025), RIVM). Developed from the start with the reworked version of the CASE software, the platform collects baseline and follow-up data while enabling centralized participant recruitment across multiple coordinated research projects. By streamlining intake and periodic follow-up engagement, the system improves long-term cohort management and enables targeted substudies. The platform also supports integration with clinical studies, as participants are invited through the portal to participate in, for instance, blood bank studies and random control trials, illustrating how CASE allows the digital research infrastructure to connect efficiently with laboratory workflows (UMC Utrecht, 2025).